Cost of Equity with and without Flotation Jarett & Sons's common stock currently trades at $22.00 a share. It is expected to pay an annual dividend of $2.75 a share at the end of the year (D1 = $2.75), and the constant growth rate is 7% a year. What is the company's cost of common equity if all of its equity comes from retained earnings? Round your answer to two decimal places. Do not round your intermediate calculations. % If the company issued new stock, it would incur a 14% flotation cost. What would be the cost of equity from new stock? Round your answer to two decimal places. Do not round your intermediate calculations. %
Answer #1

a.Cost of equity=(D1/Current price)+Growth rate

=(2.75/22)+0.07

=19.5%

b.Cost of equity=(D1/Current price(1-Flotation cost)+Growth rate

=(2.75/(22*(1-0.14))+0.07

=(2.75/18.92)+0.07

=21.53%(Approx).